Plz help m not able to prove it.To prove (Cos8•-sin8•)/(cos8•+sin8•)= tan37•

Grade:11

1 Answers

Likith Narayan ss
7 years ago
(cos 8-sin8)/(cos 8+sin8) =(cos 8(1-sin 8/cos 8))/(cos 8(1+sin8/cos8))
                                    by cancelling common cos 8 and writing tan 8=sin8/cos8
                                    =(1-tan8)/(1+tan8)
             writing 1as tan45
                                   =(tan45-tan8)/(1+tan45*tan8)
which is in the form of (tanA-tanB)/(1+tanA*tanB)=tan(A-B)
                                   =tan(45-8)
                                   =tan(37)